# Pluriclosed metrics on compact semisimple Lie groups

### v2

@@ 6. Appendix: Roots
where $\Delta\subset(\mathfrak{t}^{c})^{*}$ is the root system, i.e., $[H,E]=\alpha(H)E$ for all $H\in\mathfrak{t}^{c}$ and $E\in\mathfrak{g}_{\alpha}$. Since the Killing form $\operatorname{B}$ of $\mathfrak{g}^{c}$ is non-degenerate on $\mathfrak{t}^{c}$, for any $\alpha\in\Delta$, there is a unique $H_{\alpha}\in\mathfrak{t}^{c}$ such that $\alpha=\operatorname{B}(H_{\alpha},\cdot)$. There exist vectors $E_{\alpha}$, $\alpha\in\Delta$ such that
@@
The numbers $N_{\alpha,\beta}$ such that $[E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ satisfy the following properties:
$N_{-\alpha,-\beta}=-N_{\alpha,\beta}=N_{\beta,\alpha},\qquad N_{\alpha,\beta}=N_{\beta,\gamma}=N_{\gamma,\alpha},\qquad\mbox{if}\quad\alpha+\beta+\gamma=0,$

### v3

@@ 6. Appendix: Roots
where $\Delta\subset(\mathfrak{t}^{c})^{*}$ is the root system, i.e., $H,E]=\alpha(H)E$ for all $H\in\mathfrak{t}^{c}$ and $E\in\mathfrak{g}_{\alpha}$. Since the Killing form $\operatorname{B}$ of $\mathfrak{g}^{c}$ is non-degenerate on $\mathfrak{t}^{c}$, for any $\alpha\in\Delta$, there is a unique $H_{\alpha}\in\mathfrak{t}^{c}$ such that $\alpha=\operatorname{B}(H_{\alpha},\cdot)$. There exist vectors $E_{\alpha}$, $\alpha\in\Delta$ such that
@@
The vectors $E_{\alpha}$ can be chosen in such a way that the numbers $N_{\alpha,\beta}$ such that $E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ satisfy that (see [K, Chapter VI, Theorem 6.6])
$N_{-\alpha,-\beta}=-N_{\alpha,\beta}.$
These numbers also satisfy the following properties:
$N_{\beta,\alpha}=-N_{\alpha,\beta},\qquad N_{\alpha,\beta}=N_{\beta,\gamma}=N_{\gamma,\alpha},\quad\mbox{if}\quad\alpha+\beta+\gamma=0,$